An object is dropped from a platform 400 ft high. Ignoring wind resistance, how long will it take to reach the ground?
1
Expert's answer
2014-03-24T12:07:44-0400
s = gt^2/2
t = sqrt(2 s/g) = sqrt(2*400 ft/32 fts^-2) = 25 s